Harnett Corporation has two manufacturing departments--Molding and Assembly. The company used the following data at the beginning of the period to calculate predetermined overhead rates:

Molding Assembly Total
Estimated total machine-hours (MHs) 2,000 8,000 10,000
Estimated total fixed manufacturing overhead cost $ 8,800 $ 39,200 $ 48,000
Estimated variable manufacturing overhead cost per MH $ 1.50 $ 3.00

During the period, the company started and completed two jobs--Job E and Job M. Data concerning those two jobs follow:

Job E Job M
Direct materials $ 13,900 $ 7,800
Direct labor cost $ 21,000 $ 7,800
Molding machine-hours 1,250 750
Assembly machine-hours 1,250 6,750

Required:

a. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate that overhead rate. (Round your answer to 2 decimal places.)

b. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the amount of manufacturing overhead applied to Job E. (Do not round intermediate calculations.)

c. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours. Calculate the total manufacturing cost assigned to Job E. (Do not round intermediate calculations.)

d. Assume that the company uses a plantwide predetermined manufacturing overhead rate based on machine-hours and uses a markup of 20% on manufacturing cost to establish selling prices. Calculate the selling price for Job E. (Do not round intermediate calculations.)

e. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments. What is the departmental predetermined overhead rate in the Molding department? (Round your answer to 2 decimal places.)

f. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. What is the departmental predetermined overhead rate in the Assembly department? (Round your answer to 2 decimal places.)

g. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. How much manufacturing overhead will be applied to Job E? (Do not round intermediate calculations.)

h. Assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both production departments. Further assume that the company uses a markup of 20% on manufacturing cost to establish selling prices. Calculate the selling price for Job E. (Do not round intermediate calculations.)
Answer #1

a) Plantwide overhead rate = (2000*1.5+8000*3+48000)/10000 = 7.50 per machine hour

b) Manufacturing overhead applied to Job E = (1250+1250)*7.50 = 18750

c) Total manufacturing cost

Job E
Direct material 13900
Direct labor 21000
Overhead 18750
Total manufacturing cost 53650

d) Selling price = 53650+(53650*20%) = 64380

e) Molding department overhead rate = (8800/2000+1.50) = 5.90 per MH

f) Assembly department overhead rate = (39200/8000+3.00) = 7.90 per MH

g) Manufacturing overhead applied = (1250*5.9+1250*7.90) = 17250

h) Calculate selling price

Job E
Direct material 13900
Direct labor 21000
Overhead 17250
Total manufacturing cost 52150
Markup 10430
Selling price 62580